there are one hundred clowns how many ways can a team of 9 be created

Answer by stanbon(75887)

there are one hundred clowns how many ways can a team of 9 be created
------
100C9 = (100*99*98...*91)/(1*2*3...*9) = 1.9922...*10^12 ways
===========================
Cheers,
Stan H.
===========================